OpenStudy (konradzuse):

The scores on the Wechsler Adult Intelligence Scale are approximately Normal with μ = 100 and σ = 15. Reference: Ref 3-4 The proportion of adults with scores between 90 and 110 is closest to A. 0.250. B. 0.432. C. 0.500. D. 0.667.

OpenStudy (konradzuse):

I say D.... I'm not 100% how to solve it, I'm bguessing based on what I do know. 1 dev is 68% and 90-100 is a little less than that...

OpenStudy (konradzuse):

there is also The proportion of adults with scores above 130 is closest to A. 0.001. B. 0.023. C. 0.050. D. 0.950. How high a score is needed to be in the highest 5%? A. 115. B. 125. C. 130. D. 135. Which I say B and B

OpenStudy (konradzuse):

@EulerGroupie

OpenStudy (anonymous):

I used a z table. Starting with \[z=\frac{x-\mu}{\sigma}\] for the bottom: z=(90-100)/15=-.67 for the top: z=(110-100)/15=.67 look up the values in the z table to get the proportions below the z values I get 0.2514 and 0.7486 from the z table.